# Deep Reinforcement Learning for Bipedal Locomotion: A Brief Survey

### v1

@@ Appendix A Deep reinforcement learning algorithms
The advancement and development of RL is crucial for bipedal locomotion. Specifically, advancements in deep learning provide deep neural networks serving as function approximators to empower RL with the capability to handle tasks characterized by high-dimensional and continuous spaces, by efficiently discovering condensed, low-dimensional representations of complex data. In comparison to other robots of different morphologies, such as wheeled robots, bipedal robots feature much higher DoFs and continuously interact with environments, which results in higher requirements for the DRL algorithms. Especially in the legged locomotion field, policy gradient-based algorithms are prevalent in the field of bipedal locomotion.
Designing an effective neural network architecture is essential for tackling complex bipedal locomotion tasks. Multi-layer perceptrons (MLP), a fundamental neural network, excel in straightforward regression tasks with lower computational resource demands. A comprehensive comparison between MLP and the memory-based neural network, Long Short-Term Memory (LSTM) reveals that MLPs have an advantage in convergence speed for tasks [65]. However, LSTM, as a variant of Recurrent Neural Networks (RNN), is adept at processing data associated with time, effectively relating different states across time, and modeling key physical properties vital for periodical gaits [19] and successful sim-to-real transfer in bipedal locomotion. Additionally, Convolutional Neural Networks (CNN) specialize in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial. This diverse range of neural network architectures highlights the importance of selecting the appropriate model based on the specific requirements of the bipedal locomotion tasks.
Considering DRL alogirthms, recent bipedal locomotion studies focus on model-free reinforcement algorithms. Unlike model-based RL, which learns a model of the environment but may inherit biases from simulations that do not accurately reflect real-world conditions, model-free RL directly trains policies through environmental interaction without relying on an explicit environmental model. Although model-free RL requires more computational samples and resources, it can train a more robust policy allowing the robots to travel around challenging environments.
Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimization) and value-based approaches. Value-based methods e.g. Q-learning, Deep Q-learning (DQN) [82] only excel in discrete action space and often struggle with high dimensional action space. In contrast, policy-based methods, such as policy gradient, can handle complex tasks but are generally less sample-efficient compared to value-based methods.
@@
More advanced algorithms combine both policy-based methods and value-based methods. Actor-critic (AC) refers to a main idea simultaneously learning both a policy (actor) and a value function (critic), where it owns both advantages of both algorithms [83, 84]. Popular algorithms e.g. Trust region policy optimization (TRPO) [85] and PPO based on policy-based methods, borrow ideas from AC. Moreover, there are other novel algorithms based on the AC framework, Deep Deterministic Policy Gradient (DDPG) [86], Twin Delayed Deep Deterministic Policy Gradients (TD3) [87], A2C (Advantage Actor-Critic), and A3C (Asynchronous Advantage Actor-Critic) [88], SAC (Soft Actor-Critic) [89]. Each algorithm has its strengths considering different tasks in the bipedal locomotion scenario. There are several key factors to value these algorithms such as: sample efficiency, robustness and generalization, and implementation challenges. A comparative analysis work [90] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO that obtain relatively high score.
@@
In [91], PPO demonstrates the robustness and computational economy in complex scenarios, such as bipedal locomotion, utilizing fewer resources than TRPO. In terms of training time, PPO is much faster than SAC, and DDPG algorithms [90]. Besides, many works [19, 45, 36] have demonstrated its robustness and ease of implementation and combined with the flexibility to integrate with various neural network architectures have made PPO the most popular choice in this field. Various work has demonstrated that PPO can conduct the exploration of walking [19], jumping [37], stair climbing [38], and stepping stones [20], which demonstrates its efficiency, robustness and generalization.
@@
Additionally, the DDPG algorithm integrates the Actor-Critic framework with DQN to facilitate off-policy training, further optimizing sampling efficiency. In some explicit scenarios such as jumping, DDPG shows higher reward and better learning performance than PPO [21, 92]. TD3 is developed based on DDPG, and improve over the performance of the DDPG and SAC [89]. Soft Actor-Critic (SAC) further the agent’s exploration capabilities and sample efficiency [89]. While A2C offers improved efficiency and stability compared to A3C, the asynchronous update mechanism for A3C provides better capabilities for exploration and accelerating learning. Although these algorithms show their advancements, they are more challenging to apply due to algorithms’ complexity compared to PPO.
Appendix B Bridging sim-to-real gap
Due to the large number of interactions needed for RL algorithms, training directly on robots can lead to costly damage to hardware and the environment. Consequently, training a policy in the simulation and then deploying it to the hardware illustrates significant potential and efficiency. However, the gap between simulation and the real world remains substantial, making sim-to-real challenging. To overcome the gap, several sim-to-real approaches are developed, including dynamics randomization [36], system identification [93, 94], periodic reward composition [66], learned actuators dynamics [93, 95], regulation feedback controller [96], adversarial motion prior [18, 17].
There are two primary approaches to training policies under domain randomization. One is end-to-end training with a history of robot measurements or I/O [36] and another is policy distillation, an expert policy with environmental insights guides a student policy that learns from internal sensory feedback, such as teacher-student policy [66], RMA [64].
Details of these sim-to-real transition approaches are shown below:
1) The dynamics randomization method involves systematically varying the physical parameters of the simulated environment-such as mass, inertia, or stiffness. 2) System Identification methods develop mathematical models of dynamics from observed data, enhancing the accuracy of robots’ properties within the models, such as mass and inertia, to ensure the model faithfully represents the system’s behavior. 3) The learned actuator dynamics method utilizes experimental data from actuators to develop a model of their dynamics, achieving sim-to-real by incorporating realistic actuator behavior within the training environment. It is noticeable that the higher-level planner can also learn from reference or non-reference. 4) periodic reward composition helps capture the essential locomotion information and the periodic gait pattern is more general to adapt to uncertainty and variation in the real world. 5) The regulation feedback controller manually tunes the setting of the controller to mitigate perturbations and gaps between the sim and the real, thereby enhancing the robustness and adaptation. Key aspects of sim-to-real, including system identification, state estimation with noise measurement, and the selection of state-and-action spaces are highlighted in [34].

### v3

@@ Appendix A Deep reinforcement learning algorithms
The advancement and development of RL is crucial for bipedal locomotion. Specifically, advancements in deep learning provide deep NNs serving as function approximators to empower RL with the capability to handle tasks characterised by high-dimensional and continuous spaces, by efficiently discovering condensed, low-dimensional representations of complex data. In comparison to other robots of different morphologies, such as wheeled robots, bipedal robots feature much higher DoFs and continuously interact with environments, which results in higher requirements for the DRL algorithms. Especially in the legged locomotion field, policy gradient-based algorithms are prevalent in the field of bipedal locomotion.
Designing an effective NN architecture is essential for tackling complex bipedal locomotion tasks. Multi-layer perceptrons (MLP), a fundamental NN, excel in straightforward regression tasks with lower computational resource demands. A comprehensive comparison between MLP and the memory-based NN, LSTM reveals that MLPs have an advantage in convergence speed for tasks [77]. However, LSTM, as a variant of Recurrent Neural Networks (RNN), is adept at processing data associated with time, effectively relating different states across time, and modelling key physical properties vital for periodical gaits [28] and successful sim-to-real transfer in bipedal locomotion. Additionally, Convolutional Neural Networks (CNN) specialise in spatial data processing, particularly for image-related tasks, making them highly suitable for environments where visual perception is crucial. This diverse range of NN architectures highlights the importance of selecting the appropriate model based on the specific requirements of the bipedal locomotion tasks.
Considering DRL algorithms, recent bipedal locomotion studies focus on model-free RL algorithms. Unlike model-based RL, which learns a model of the environment but may inherit biases from simulations that do not accurately reflect real-world conditions, model-free RL directly trains policies through environmental interaction without relying on an explicit environmental model. Although model-free RL requires more computational samples and resources, it can train a more robust policy allowing the robots to travel around challenging environments.
Many sophisticated model-free RL algorithms exist, which can be broadly classified into two categories: policy-based (or policy optimisation) and value-based approaches. Value-based methods e.g. Q-learning, State-Action-Reward-State-Action (SARSA), Deep Q-learning (DQN) [99] only excel in discrete action space and often struggle with high dimensional action space. Q-learning is an off-policy algorithm that directly learns the optimal Q-values, allowing it to derive the best possible actions irrespective of the current policy. SARSA, an on-policy variant, updates its Q-values based on the actual actions taken, making it robust in environments where the policy evolves during learning. DQN extends Q-learning by using deep NNs to approximate Q-values, enabling the algorithm to tackle complex state spaces, though it still faces challenges with high-dimensional action spaces due to the difficulty of accurate value estimation. In contrast, policy-based methods, such as policy gradient, can handle complex tasks but are generally less sample-efficient compared to value-based methods.
@@
More advanced algorithms combine both policy-based methods and value-based methods. Actor-critic (AC) refers to a main idea simultaneously learning both a policy (actor) and a value function (critic), where it owns both advantages of both algorithms [100, 101]. Popular algorithms e.g. Trust region policy optimisation (TRPO) [102] and PPO based on policy-based methods, borrow ideas from AC. Moreover, there are other novel algorithms based on the AC framework, Deep Deterministic Policy Gradient (DDPG) [103], Twin Delayed Deep Deterministic Policy Gradients (TD3) [104], A2C (Advantage Actor-Critic), and A3C (Asynchronous Advantage Actor-Critic) [105], SAC (Soft Actor-Critic) [106]. Each algorithm has its strengths considering different tasks in the bipedal locomotion scenario. There are several key factors to value these algorithms such as: sample efficiency, robustness and generalisation, and implementation challenges. A comparative analysis work [107] illustrates that SAC-based algorithms excel in stability and achieve the highest scores, while their training efficiency significantly trails behind that of PPO that obtain relatively high score.
@@
In [108], PPO demonstrates the robustness and computational economy in complex scenarios, such as bipedal locomotion, utilising fewer resources than TRPO. In terms of training time, PPO is much faster than SAC, and DDPG algorithms [107]. Besides, many works [28, 55, 44] have demonstrated its robustness and ease of implementation and combined with the flexibility to integrate with various NN architectures have made PPO the most popular choice in this field. Various work has demonstrated that PPO can conduct the exploration of walking [28], jumping [45], stair climbing [46], and stepping stones [29], which demonstrates its efficiency, robustness and generalisation.
@@
Additionally, the DDPG algorithm integrates the Actor-Critic framework with DQN to facilitate off-policy training, further optimising sampling efficiency. In some explicit scenarios such as jumping, DDPG shows higher reward and better learning performance than PPO [30, 109]. TD3 is developed based on DDPG, and improve over the performance of the DDPG and SAC [106].
SAC improves exploration through its stochastic policy and entropy-regularised objective, which encourages the agent to maintain randomness in its actions, balancing exploration and exploitation more effectively than DDPG and TD3. Unlike PPO, which is an on-policy algorithm, SAC’s off-policy nature allows it to leverage a replay buffer, reusing past experiences to train without requiring constant interactions with the environment. This, combined with entropy maximisation, enables SAC to achieve faster convergence in complex environments where exploration is essential. SAC is also known for its stability and strong performance across a wide range of tasks [106]. While A2C offers improved efficiency and stability compared to A3C, the asynchronous update mechanism for A3C provides better capabilities for exploration and accelerating learning. Although these algorithms show their advancements, they are more challenging to apply due to algorithms’ complexity compared to PPO.
Appendix B Training Simulation Environment
The development of DRL algorithms and sim-to-real techniques has shown the requirement for high-quality simulators. Creating a reliable simulation environment and conducting RL training is difficult. Through the literature, several simulators are available, including Isaac Gym [72], robocup3D [73], Openai Gym [74], MuJoCo [75].
OpenAI developed Gym and Gymnasium providing a simplified environment for fast testing RL algorithms with the simplified bipedal model. Also, RoboCup serves as a platform for RL research and development. On the other hand, MuJoCo and Gazebo developed by DeepMind respectively are comprehensive physics engines that support robotics research and development, along with other applications needing physics simulation. Isaac Gym, a cutting-edge physics simulation and RL framework developed by NVIDIA, is a high-performance platform for training agents in intricate and lifelike environments.
One of the most crucial things is the parallelisation strategy and GPU simulations. For instance, Isaac Gym was developed to maximise the throughput of physics-based machine learning algorithms with particular emphasis on simulations that require large numbers of environment instances executing in parallel. Running the physics simulation on GPU can result in significant speedups, especially for large scenes with thousands of individual actors.